OpenStudy (anonymous):

What is the slope of the line that passes through the pair of points(-5/3) and (-2, 9/)?

OpenStudy (luigi0210):

Use the slope formula: \[\LARGE m=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]

OpenStudy (luigi0210):

Then after the slope, use the point-slope formula: \[\LARGE y-y_{1}=m(x-x_{1})\]

OpenStudy (mathmale):

Hey, Frinchie, if you meant the points (-5,3) and (-2,9), you could label \[x _{1}=-5,y _{1}=3,\] and so on. Doesn't matter which of the two points you think of as #1.

OpenStudy (anonymous):

I can't figure out the slope formula I can't find the way to subtract the fractions from the other number(s)

OpenStudy (mathmale):

OK. First of all, our friend Luigi is right on target in providing the slope formula: \[m=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (mathmale):

If, as I suggested earlier, \[x _{1}=-5,y _{1}=3,x _{2}=-2,andy _{2}=9,\]

OpenStudy (mathmale):

all you need to do next is to substitute those values into Luigi's slope formula. Would you please try doing that?

OpenStudy (anonymous):

so, 6 over -7???

OpenStudy (mathmale):

Please double check your work. I got 6 over 3.

OpenStudy (mathmale):

Please pay special attention when you subtract negative numbers.

OpenStudy (mathmale):

We agree on the 6 in the numerator. As for the denominator, since \[x _{2}=-2 and x _{1}=-5, we need (-2)-(-5).\] What does that come to?

OpenStudy (anonymous):

I re-did my work and yeah it =6 over 3, so that would mean that we need to grab the next equation that Luigi gave us?

OpenStudy (mathmale):

Yes. But first, please reduce 6/3. That's your slope. In simplest form, your slope has what value?

OpenStudy (anonymous):

So I reduced it to 1 over 2, or 1/2

OpenStudy (mathmale):

As I see it, Frinchie, 6/3 reduces to 2/1. What do you think?

OpenStudy (mathmale):

\[\frac{ 6 }{ 3 }=\frac{ 2*3 }{ 1*3 }=?\]

OpenStudy (anonymous):

That looks right,

OpenStudy (mathmale):

OK. Luigi's second equation is the point-slope formula for a straight line, used when we know one point on the line in question and also know the slope of the line. We're in that situation now. You may pick EITHER of the 2 given points, as well as use the slope m=2 that we just got. Which point would you like to use? Really, it doesn't matter which one you choose.

OpenStudy (anonymous):

Will 3 work?

OpenStudy (mathmale):

OK. Here's that formula: \[y-y _{1}=m(x-x _{1}).\]

OpenStudy (mathmale):

We found that the slope, m, is 2, didn't we?

OpenStudy (mathmale):

Since you mention "3" I'm assuming that you chose to use the first point: (-5,3). That's just fine. You could have used the other point and still arrived at the correct equation of the line in question.

OpenStudy (mathmale):

So, Frinchie, your slope is m=2, your \[x _{1}=-5,y _{1}=3.\]

OpenStudy (mathmale):

Go ahead and substitute those values into Luigi's "point-slope form".

OpenStudy (anonymous):

So we currently have Y-3=2(X+5) as are equation?

OpenStudy (mathmale):

Yes, that's perfect. Excellent! Let's summarize what you've done. You've taken the two given points and from them you've found the slope of the line, which is 2.

OpenStudy (mathmale):

You then substituted the appropriate coordinates from the first point into that point-slope equation, along with the slope m=2. You've ended up with the equation of the line in question. You could either stop there, or

OpenStudy (mathmale):

you could add 3 to both sides of the equation to solve it for y.

OpenStudy (anonymous):

We can stop there, can I ask one more question?

OpenStudy (mathmale):

Of course, please go ahead!

OpenStudy (anonymous):

Write y=(-3 over 4)x=3

OpenStudy (anonymous):

x+3*

OpenStudy (mathmale):

Let me give you my interpretation: Did you mean:\[y=\frac{ 3 }{ 4 }x+3?\]

OpenStudy (mathmale):

Oops..I missed the negative sign.

OpenStudy (anonymous):

yup

OpenStudy (mathmale):

What aspect of this equation would you like to discuss?

OpenStudy (anonymous):

How to do this I havn't really understood this whole concept

OpenStudy (mathmale):

Mind saying what it is that you would like to "do"? What you have there is just the equation of a straight line in slope-intercept form.

OpenStudy (mathmale):

Note that this form is different from the point-slope form Luigi and we used before.

OpenStudy (anonymous):

Nevermind I figured it out thanks for the help!!!

OpenStudy (mathmale):

Very happy to work with you. Encourage you to come back to OpenStudy with further questions. Just out of curiosity, where are you located? I'm in S. California.

OpenStudy (anonymous):

Alpine Utah

OpenStudy (mathmale):

Cool. What a beautiful state! All the best to you, and Happy New Year!

OpenStudy (anonymous):

Happy new year to you aswell

OpenStudy (mathmale):

So long for now! over and out.